In a right traingle fond the length of the side not given
b=1 and c=sqt26
thank you

Answer by Alan3354(69443)

In a right traingle fond the length of the side not given
b=1 and c=sqt26
------------
a^2 + b^2 = c^2
a^2 + 1 = 26
a = 5